# How Exclusive are Ethereum Transactions?

### v1

@@ Abstract
We analyze 15,097 blocks proposed for inclusion in Ethereum’s blockchain over an 8-minute window on December 3, 2024, during which 38 blocks were added to the chain. We classify transactions as exclusive—present only in blocks from a single builder—or private—absent from the public mempool but included in blocks from multiple builders. We find that exclusive transactions account for 84% of the total fees paid by transactions in winning blocks. Furthermore, we show that exclusivity cannot be fully explained by exclusive relationships between senders and builders: about 7% of all exclusive transactions included on-chain, by value, come from senders who route exclusively to a single builder. Analyzing transaction logs shows that some exclusive transactions are duplicates or variations of the same strategy, but even accounting for that, the share of the total fees paid by transactions in winning blocks is at least 77.2%. Taken together, our findings highlight that exclusive transactions are the dominant source of builder revenues.
@@ 1 Introduction
Our paper addresses this challenge using a novel dataset that covers about 28% of all blocks proposed for inclusion in Ethereum during an 8-minute window on December 3, 2024, corresponding to 39 new blocks. [^2] The dataset contains all blocks (winning and non-winning) submitted via the “Agnostic relay” during the study period. They are therefore not a random sample of all proposed blocks (see Section 3 for details and limitations). To our knowledge, together with the companion paper Canidio and Pahari (2025), this is the first study to analyze the content of non-winning Ethereum blocks. It includes 15,097 proposed blocks, 10,793 unique transactions (identified by their hash), and 2,380,014 transaction–block pairings. By comparing the content of proposed blocks with that of the public mempool (sourced from https://mempool.guru/), we classify each transaction as:
@@
Of the 5,576 transactions in our dataset that were ultimately included on-chain, 4,124 were public, 810 private, and 642 exclusive. Exclusive transactions, though fewer in number, are by far the most valuable: they account for 84% of all fees paid in winning blocks. This result confirms their central role in block building. We also find that almost all (i.e., 91%) transactions that never make it on-chain are exclusive.
We then study whether transactions change status over time. Among the transactions that first appear as exclusive, 162 remain exclusive during multiple bidding cycles, 174 transition to private in a subsequent bidding cycle, and 12 exclusive transactions later appear in the public mempool. These 12 transactions seem to be user-generated, as they swap on DEX pools using well-known routers (0x and 1inch). Their initial exclusivity is puzzling: ordinary users typically gain no benefit from sending a transaction to a single builder rather than sharing it privately via a known private mempool or broadcasting it publicly. More surprisingly, these transactions later appear in the public mempool, meaning they were first delayed and then exposed to attacks. We speculate that a private mempool operator may have shared them exclusively with one builder, falling back to the public mempool if that builder failed to win. [^3] The presence of exclusive users’ transactions is notable because four of the five major private mempool operators explicitly state that they share transactions with multiple builders (MEV Blocker, Flashbots Protect, Blink, and Merkel; see Janicot and Vinyas, 2025). In contrast, the fifth operator, Metamask Smart Transactions, does not publicly disclose its sharing policies.
We also find cases of transactions changing status within a bidding cycle. We identify 19 cases where transactions start as exclusive to a single builder during the initial part of the auction cycle, but later in the same cycle appear in blocks from other builders. Most of these were submitted via Flashbots Protect: they were initially exclusive to the Flashbots builder before being propagated more broadly.
@@
To investigate the origin of these exclusive transactions, we analyze the relationship between transaction senders and builders. When we simply count transactions, we find that the vast majority of exclusive transactions can be attributed to senders interacting exclusively with one builder. However, when we focus on value and only consider transactions that are included on-chain, according to our preferred operationalization of “transaction sender” only about 7% of value can be attributed to exclusive relationships between searchers and builders, which therefore do not seem very important in explaining the overall importance of exclusive transactions.
We conclude by examining cases where the same sender submits exclusive transactions to multiple builders within a single auction cycle. Using transaction logs, we identify three patterns: (i) 171 transactions have different hashes but are functionally identical to another exclusive transaction in the same cycle, executing the same operation with the same fee; (ii) 494 transactions are identical in execution to another exclusive transaction in the same cycle, but differ in the fee attached; and (iii) 1,179 transactions are similar but not identical to another exclusive transaction in the same cycle, swapping in the same DEX pools but different amounts. While analyzing logs refines our classification of exclusivity, the number of affected transactions is small and does not change our main conclusion.
@@
Finally, in our companion paper Canidio and Pahari (2025), we use the same dataset to address two related questions: whether block selection matters for users, and competition between arbitrageurs. The paper shows that many user transactions are delayed: although proposed during a bidding cycle, they are not included in the winning block. Furthermore, they also identify two arbitrage bots trading between decentralized (DEX) and centralized (CEX) exchanges. Analyzing the bidding dynamics, they estimate that the implied CEX prices at which they trade USDC/WETH and USDT/WETH are 3.4–4.2 basis points better than the contemporaneous Binance price.
@@ 2 Background on PBS
As noted above, in Ethereum, the proposer almost always delegates block construction to builders. Builders collect both user transactions and those submitted by searchers, automated bots that generate transactions in response to market conditions (e.g., price fluctuations) or to other transactions pending in the public mempool. Searchers play a central role in decentralized finance: they engage in arbitrage across DEXs and CEXs, liquidate under-collateralized loans, or exploit temporary price discrepancies. Because their strategies often depend on speed and precise ordering, searchers typically route transactions to builders via private channels. Users can submit transactions in different ways. The default is to broadcast them to the public mempool, where all builders can observe them. This, however, exposes users to attacks. [^4] For example, if a user swaps assets on an Automated Market Maker (AMM), a malicious searcher can mount a “sandwich” attack: front-running the swap with the same trade and then back-running it with the opposite trade. The attacker buys low and sells high, forcing the victim to execute at worse prices. To reduce this risk, many users rely on private mempools, publicly available services which distribute transactions privately to a subset of builders.
@@ 3 Dataset
  • Hashes of blocks submitted to all major relays: Flashbots, BloXroute, Manifold, Eden, Ultra Sound, SecureRpc, and Aestus. Because any difference in content or bid produces a unique hash, this lets us check whether blocks in our dataset were also sent elsewhere. We find that our 15,097 blocks represent 28.3% of all blocks submitted to major relays during the study period.
  • Data on transactions from the public mempool and from two private mempools operators (MEV Blocker and Flashbots Protect), obtained via https://mempool.guru/ and https://Dune.com.
@@
An important detail of PBS for our analysis is that a relay does not run an auction for a given slot if the assigned proposer is not registered with it. This occurs in our data: of the 39 slots we study, 11 show no bidding activity because the proposers were not connected to the Agnostic relay. [^7] These are the slots leading to blocks 21,322,624; 21,322,625; 21,322,627; 21,322,629; 21,322,633; 21,322,634; 21,322,636; 21,322,644; 21,322,646; 21,322,647; and 21,322,651. Also, each block in our dataset has two timestamps: received at (i.e., received from the builder) and made available at (i.e., made available to the proposer). The gap between them reflects the time needed for the relay to simulate the block and verify its validity. [^8] Some blocks are treated “optimistically”—they are made available immediately, with simulation performed in the background. If the simulation later fails, the builder’s next submission is not treated optimistically. Such blocks account for only 3.8% of our sample. This simulation delay is non-trivial: the median is 0.76 seconds, with the 75th percentile at 1.5 seconds. In what follows, we use the received at timestamp, as it better captures the events we study—such as when a builder submits a block or when a transaction first appears.
As noted earlier, our primary dataset covers 15,097 blocks submitted across 28 bidding cycles. The number of blocks per slot ranges from 220 to 951, with a mean of 539 and a median of 509. We can attribute 14,043 of these blocks to 23 known builders. The most active are Titan Builder (7,024 blocks, 46.5%), Rsync Builder (2,259, 15.6%), Flashbots (1,936, 12.8%), and Beaverbuild (1,418, 9.4%). Most blocks in our dataset—12,895 in total—were also submitted to at least one other relay. However, when comparing across all relays, we find an imbalance. Titan Builder submitted 13,679 blocks overall (25.6%), Rsync Builder 7,936 (14.9%), Flashbots 2,626 (4.9%), and Beaverbuild 21,447 (40.2%). Thus, Beaverbuild contributes far fewer blocks to Agnostic Relay than to other relays. We recognize this imbalance as a limitation of our dataset.
@@ 4 Results
4.1 Transactions’ transitions.
Among the 1,330 transactions observed across multiple bidding cycles, public transactions remain public in all subsequent cycles. Only a handful of private transactions (13 out of 208) later appear in the public mempool. By contrast, the majority of exclusive transactions transition to private or even public. Of the transactions that begin as exclusive, 162 remain exclusive across all observed cycles, while 174 transition to private. [^9] We also find 33 transactions that were initially exclusive to a builder and later included in a subsequent cycle won by the same builder. Because those cycles are not in our primary dataset, we cannot determine whether the transactions remained exclusive or became private. In addition, 12 transactions that were initially exclusive eventually surfaced in the public mempool. Nine of these were present in two cycles and three in three cycles. All appear to be user-generated: three used the 0x router and four the 1inch router. [^10] These transactions are: 0x1daea1584d684385fb25209bad3a49c54a1e2c500e27543334c7b3e695a12ffb, 0x6605c1c36fcd3d696844519d6ff7199fe4ddd34fb4f3e0ed4da71aedfcf1c633, 0x3dadbab8e85b6e48a2c33f9029ea4ebebdd3def8b9d907faff15ef81527de968, 0x9c7364eddb94fe659bc8104585ff86a8a234480fe00a9f9aa3eb410e572ebdf2, 0x9d476dbc524651cbe26bcafab99f6fe40a823a5588fa5cd5002e5b414a8a2f8c, 0x417440b083bdcb529663c6615c32d89a92beb7863f1a5d6b12bca5b94f4a0f87, 0xe56caee27e79dee8a4c7ed6f945ccbc2cc733d761e19bd68329dc364c61e73a7, 0xae5b88bc0bffa04811eda0dfdcd7224599e2757339e18b2d57f4f40d61902a4c, 0xd41139b67a193045b4d60266c084f6dadd0cb95dccd7b2a06ea68653d6a313ec, 0x597fc0cb0610d1e3ea2baf0f6f90d0d741a6429bfdd9b375bfd2493b70d41dcd, 0x44e8105d5c672195877dc1ed970511e0f0a15fc25cbcafe71d51d428d54c215d, 0xd74718e6d8d58bc33334c0c2f6642d9dec49496329329d8fdcb1e698e81abc94. The existence of these transactions is problematic, as they were first delayed and later exposed in the public mempool, leaving users vulnerable to attacks.
Of Titan’s 17 cases, 8 came through the Flashbots private mempool. Five originated from a single searcher address1111110x89a99a0a17d37419f99cf8dc5ffa578f3cdb58b5 and appear in Flashbots Protect’s dataset as shared with more than 15 builders. These transactions remained exclusive to Flashbots Builder for about 2.6 seconds before entering Titan’s blocks. Another five transactions from the same address follow the same pattern (initial exclusivity to Flashbots, later appearance in Titan’s blocks) but were never included on-chain. It is plausible that they were also shared via Flashbots Protect, which records only transactions that are eventually included.
In summary, transaction sharing between builders during the bidding cycle does occur but is relatively rare. When it happens, it typically involves transactions sent via Flashbots Protect that are initially exclusive to Flashbots Builder.
@@ 4.2 Relationship between searchers and builders
To better interpret these results, we ask whether exclusive transactions are the outcome of exclusive relationships between senders and builders, or, instead, the same sender submits different exclusive transactions to different builders. In the latter case, the natural question is how these transactions differ across builders.
@@
To study the relationships between the senders of the 5,228 exclusive transactions and builders, we must first define “sender.” One option is the from address. However, many searchers use multiple from addresses while embedding their logic in smart contracts that automatically generate transactions. For searchers, therefore, a more reliable identifier is the to address, corresponding to the address of the smart contract used by the searcher. We may also want to restrict attention to senders with a minimum number of transactions. Otherwise, a sender with only one or two exclusive transactions will mechanically appear to work exclusively with a single builder, which may reflect limited data rather than a genuine relationship.
Tables 1 and 2 group transactions by the number of builders each sender interacts with, under different definitions of “sender.” Tables 1 reports the resulting transaction counts and the total fees paid by those transactions, while Table 2 reports the same but only for transaction that are ultimately included on chain. While transaction counts vary somewhat with the definition of “sender,” the fee data in Table 1 are more robust: by value, roughly half of all exclusive transactions arise from exclusive relationships between senders and builders. However, once we restrict to on-chain transactions, this share drops sharply. Using to addresses—which better identify searchers—only around 7% of the value of exclusive transactions that land on-chain can be traced to exclusive relationships. Taken together, these results suggest that exclusive relationships between senders and builders play only a limited role in explaining why, as noted earlier, 85% of the value of on-chain transactions comes from exclusives.

### v2

@@ Abstract
We analyze 15,097 blocks proposed for inclusion in Ethereum’s blockchain over an eight-minute window on December 3, 2024, during which 38 blocks were added to the chain. We classify transactions as exclusive—appearing only in blocks from a single builder—or private — absent from the public mempool but included in blocks from multiple builders. We find that exclusive transactions account for between 77.2% and 84% of the total fees paid by transactions in winning blocks. Moreover, we show that exclusivity cannot be fully attributed to persistent relationships between senders and builders: only about 7% of all on-chain exclusive transaction value originates from senders who route exclusively to one builder. Finally, we observe that transaction exclusivity is dynamic. Some transactions are exclusive at the start of a bidding cycle but later appear in blocks from multiple builders. Other transactions remain exclusive to a losing builder for two or three cycles before appearing in the public mempool. These transactions are therefore delayed and then exposed to potential attacks.
@@ 1 Introduction
Our paper addresses this challenge using a novel dataset containing 15,097 blocks proposed for inclusion in Ethereum during an eight-minute window on December 3, 2024, corresponding to 39 new blocks. [^2] The dataset includes all blocks (winning and non-winning) submitted via the “Agnostic relay” during the study period. They are therefore not a random sample of all proposed blocks (see Section 3 for details and limitations). To our knowledge, together with the companion paper Canidio and Pahari (2025), this is the first study to analyze the content of non-winning Ethereum blocks. The dataset contains 10,793 unique transactions (identified by their hash) and 2,380,014 transaction–block pairings. By comparing the content of proposed blocks with that of the public mempool (sourced from https://mempool.guru/), we classify each transaction as:
@@
We apply this classification both globally and within shorter time intervals. Applying the classification over shorter intervals allows us to track transitions between statuses over time. For instance, a transaction may be exclusive initially but later become public or private.
@@
Of the 5,576 transactions in our dataset that were ultimately included on-chain, 4,124 were public, 810 private, and 642 exclusive. Exclusive transactions, though fewer in number, are by far the most valuable: they account for 84% of all fees paid by transactions included in winning blocks. This result confirms their central role in block building. We also find that almost all (i.e., 91%) transactions that are never included in a winning block are exclusive.
Because we identify transactions by their hash, it is possible that two exclusive transactions available to different builders perform identical or nearly identical operations, meaning they may functionally be the same transaction and thus are not truly exclusive. To examine this possibility, we study cases in which the same sender submits exclusive transactions to multiple builders within a single auction cycle. We then use their transaction logs to identify cases in which the same sender sends exclusive transaction that are:
  • functionally identical to each other, executing the same operation with the same fee;
  • identical in execution but different in the attached fee; and
  • similar but not identical, performing swaps in the same DEX pools but with different amounts.
While this log-based analysis refines our classification of exclusivity, the number of affected transactions is small and does not alter our main conclusion: even if all such transactions were reclassified from “exclusive” to “private,” exclusive transactions would still account for a large share (77.2%) of total fees paid by transactions included in winning blocks.
@@
To investigate the origin of exclusive transactions, we analyze the relationship between transaction senders and builders. Counting transactions alone, we find that the vast majority of exclusive transactions can be attributed to senders that interact exclusively with a single builder. However, when we shift focus from transaction counts to transaction value and consider only those that are ultimately included on-chain, the picture changes. Under our preferred operational definition of a “transaction sender,” only about 7% of the total value of exclusive transactions can be attributed to exclusive relationships between searchers and builders. Such relationships therefore appear to play a limited role in explaining the overall value of exclusive transactions.
Finally, we study whether transactions change status over time, starting with changes across bidding cycles. Among transactions that first appear as exclusive, 162 remain exclusive over multiple cycles, 174 transition to private in a subsequent cycle, and 12 later appear in the public mempool. These 12 transactions appear to be user-generated, as they involve swaps on DEX pools via well-known routers such as 0x and 1inch. Their initial exclusivity is puzzling, as ordinary users typically gain no benefit from sending transactions directly to a single builder instead of submitting them to a private relay or the public mempool. Even more surprisingly, these transactions eventually surface in the public mempool, indicating that they were initially delayed and later exposed to potential attacks (one was in fact the victim of a sandwich attack). In four cases, the same builder who initially had exclusive access ultimately included the transaction on-chain. We speculate that a private mempool operator may have shared these transactions exclusively with one builder, reverting to the public mempool when that builder failed to win. [^3] Private mempools are services that allow users to send transactions directly to one or more builders without broadcasting them to the public mempool. At present, there are five major private mempool operators. Four of them—MEV Blocker, Flashbots Protect, Merkel, and Blink—state in their documentation that they share transactions privately with multiple builders (on the differences between those operators, see Janicot and Vinyas, 2025). Instead, the fifth operator, Metamask Smart Transaction, does not disclose its sharing policy.
We also observe transactions changing status within a single bidding cycle. Specifically, we identify 19 cases in which transactions are initially exclusive to one builder at the start of the auction cycle but later, within the same cycle, appear in blocks proposed by other builders. Most of these transactions were submitted via Flashbots Protect: they were initially exclusive to the Flashbots builder and only later propagated more broadly. We speculate that the Flashbots builder may temporarily retain transactions submitted through Flashbots Protect as exclusive, releasing them to other builders once it becomes apparent that it is unlikely to win the block.
@@
Finally, in our companion paper Canidio and Pahari (2025), we use the same dataset to address two related questions: whether block selection matters for users, and competition between arbitrageurs. We shows that many user transactions are delayed: although proposed during a bidding cycle, they are not included in the winning block. Furthermore, we identify two arbitrage bots trading between decentralized (DEX) and centralized (CEX) exchanges. Analyzing the bidding dynamics, they estimate that the implied CEX prices at which they trade USDC/WETH and USDT/WETH is 2.8 basis points better than the contemporaneous Binance price.
@@ 2 Background on PBS
As noted above, in Ethereum, the proposer almost always delegates block construction to builders. Builders collect both user transactions and those submitted by searchers—automated bots that generate transactions in response to market conditions (e.g., price fluctuations) or to other transactions pending in the public mempool. Searchers play a central role in decentralized finance: they engage in arbitrage across DEXs and CEXs, liquidate under-collateralized loans, or exploit temporary price discrepancies between DEXs. Because their strategies often depend on speed and precise ordering, searchers typically route transactions to builders via private channels.
Users can submit transactions in different ways. The default is to broadcast them to the public mempool, where all builders can observe them. This, however, exposes users to attacks. [^4] For example, if a user swaps assets on an Automated Market Maker (AMM), a malicious searcher can mount a “sandwich” attack: front-running the swap with the same trade and then back-running it with the opposite trade. The attacker buys low and sells high, forcing the victim to execute at worse prices. To reduce this risk, many users rely on private mempools, publicly available services which distribute transactions privately to a subset of builders.
@@ 3 Dataset
  • Hashes of blocks submitted to all major relays: Flashbots, BloXroute, Manifold, Eden, Ultra Sound, SecureRpc, and Aestus. Because any difference in content or bid produces a unique hash, this lets us check whether blocks in our dataset were also sent elsewhere.
  • Data on transactions submitted via the public mempool, obtained from https://mempool.guru/.
  • Data on transactions submitted via two private mempools, MEV Blocker and Flashbots Protect, obtained from https://Dune.com.
@@
An important detail of PBS for our analysis is that a relay does not run an auction for a given slot if the assigned proposer is not registered with it. This occurs in our data: of the 39 slots we study, 11 show no bidding activity because the proposers were not connected to the Agnostic relay. [^7] These are the slots leading to blocks 21,322,624; 21,322,625; 21,322,627; 21,322,629; 21,322,633; 21,322,634; 21,322,636; 21,322,644; 21,322,646; 21,322,647; and 21,322,651. By comparing the hash of blocks in our primary dataset with those submitted to the other major relays, we find that our 15,097 blocks represent 28.3% of all blocks submitted to major relays during the 28 auction cycles for which we have data.
Also, each block in our dataset has two timestamps: received at (i.e., received from the builder) and made available at (i.e., made available to the proposer). The gap between them reflects the time needed for the relay to simulate the block and verify its validity. [^8] Some blocks are treated “optimistically”—they are made available immediately, with simulation performed in the background. If the simulation later fails, the builder’s next submission is not treated optimistically. Such blocks account for only 3.8% of our sample. This simulation delay is non-trivial: the median is 0.76 seconds, with the 75th percentile at 1.5 seconds. In what follows, we use the received at timestamp, as it better captures the events we study—such as when a builder submits a block or when a transaction first appears.
As noted earlier, our primary dataset covers 15,097 blocks submitted across 28 bidding cycles. Across those cycles, the number of blocks per slot ranges from 220 to 951, with a mean of 539 and a median of 509. We can attribute 14,043 of these blocks to 23 known builders. The most active are Titan Builder (7,024 blocks, 46.5%), Rsync Builder (2,259, 15.6%), Flashbots (1,936, 12.8%), and Beaverbuild (1,418, 9.4%). Most blocks in our dataset—12,895 in total—were also submitted to at least one other relay. However, when comparing across all relays, we find an imbalance. Titan Builder submitted 13,679 blocks overall (25.6%), Rsync Builder 7,936 (14.9%), Flashbots 2,626 (4.9%), and Beaverbuild 21,447 (40.2%). Thus, Beaverbuild contributes far fewer blocks to Agnostic Relay than to other relays. We recognize this imbalance as a limitation of our dataset.
@@ 4 Results
4.1 Logs-based classification
The above method treats transactions with different hash as different transactions. However, a transaction may have a different hash from another while performing the same operations or similar ones. As a result, our method may overestimate the share of exclusive transactions. To address this concern, in this section we refine our approach by using transaction logs, that is, the sequence of operations and corresponding outputs generated by simulating the blocks that included those transactions.
The method above treats transactions with different hashes as distinct. However, two transactions may have different hashes while performing the same or very similar operations. As a result, our method may overestimate the share of exclusive transactions. To address this concern, we refine our approach in this section by comparing transaction logs, that is, the sequence of operations and corresponding outputs generated by simulating the blocks that included the transactions.
We identify exclusive transactions originating from the same sender but submitted to different builders within the same auction cycle. To do so, we must first define what constitutes a "sender." A natural option is the transaction’s from address. However, many searchers use multiple from addresses while embedding their logic in smart contracts that automatically generate transactions. For this reason, the to address—that is, the address of the smart contract called by the transaction—is a more reliable identifier of the sender, particularly in the case of searchers. Since multiple from addresses can call the same contract, grouping transactions by to address increases the likelihood of detecting functionally similar transactions that differ only in hash. In this sense, using to addresses offers a stricter test of our hash-based classification of transactions as exclusive, private, or public. We therefore adopt the to address as the definition of sender in our analysis.
We we simulate these transactions to extract their logs, we identify 665 such transactions producing 279 unique logs. Nearly all of them (630 transactions) are swaps in DEX pools, primarily on Uniswap V2, Uniswap V3, Sushiswap, Pancakeswap, or Curve. We categorize these transactions in two groups:
We find 171 partially exclusive transactions (59 unique logs) with a degree of exclusivity equal to zero—that is, the sender submitted the same transaction (same log, same fee) to at least two builders. We found another 494 partially exclusive transactions (220 unique logs) with different fees across builders. Among these, the average degree of exclusivity is 0.004178 ETH (or 23% of the value of the second-highest fee). We also identify 1,179 “similar” exclusive transactions grouped into 162 log sets (861 unique logs), where transactions differ only in the swap amount. Comparing the fees attached to these transactions, we find that the average gap between the highest payment to one builder and the second-highest payment to another builder is almost 70% of the second-highest payment. This suggests that the differences between these transactions are economically meaningful.
The takeaway is that not all transactions that appear “exclusive” at the hash level remain exclusive once we examine their logs: some are duplicates submitted to multiple builders, while others are variations of the same strategy. Quantitatively, however, these cases are limited. Also, if we change the classification of partially exclusive and similar transactions to “private”, we find that 77.2% of fees paid by transactions that are included on-chain are from exclusive transactions (vs 84% with our initial classification). The main result, therefore, stands: exclusive transactions remain the key driver of builders’ profitability.
@@ 4.2 Relationship between searchers and builders
To better interpret these results, we study whether exclusive transactions are the outcome of exclusive relationships between senders and builders, or, instead, the same sender submits different exclusive transactions to different builders.
@@
Tables 1 and 2 group transactions by the number of builders each sender interacts with, using both the to and the from definitions of “sender” (see the previous subsection) and with and without filtering out senders who send fewer than five transactions. [^10] The rationale behind the filtering is that a sender with only one or two exclusive transactions will mechanically appear to work exclusively with a single builder, which may reflect limited data rather than a genuine relationship. Table 1 reports the resulting transaction counts and the total fees paid by those transactions, while Table 2 reports the same but only for transaction that are ultimately included on chain.
Tables 1 and 2 group transactions by the number of builders each sender interacts with. We use both definitions of “sender” discussed in the previous subsection—to and from addresses—and present results both with and without filtering out senders who submit fewer than five transactions. [^11] This filter helps address a possible mechanical biases: a sender with only few transactions is more likely to appear to interact with just one builder, which may reflect data sparsity rather than a true exclusive relationship. Table 1 reports transaction counts and the total fees paid, while Table 2 restricts the analysis to transactions that were ultimately included on-chain.
In Table 1 transaction counts vary somewhat with the definition of “sender” but the fee data are more robust: by value, roughly half of all exclusive transactions arise from exclusive relationships between senders and builders. However, once we restrict to on-chain transactions, this share drops sharply. Table 2 shows that if we use to addresses—which better identify searchers—then only around 7% of the value of exclusive transactions that land on-chain can be traced to exclusive relationships. Taken together, these results suggest that exclusive relationships between senders and builders play only a limited role in explaining why, as noted earlier, 85% of the value of on-chain transactions comes from exclusive transactions.